OpenStudy (anonymous):

Find first and second derivative, the intervals of which the function is increasing,decreasing,concave up, and concave down, and locate any relative maxima,minima

OpenStudy (anonymous):

\[f(x)=4x^3-6x^2-72x+20\]

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

please help me

zepdrix (zepdrix):

Were you able to find the first derivative? :o

OpenStudy (anonymous):

i think its 12x^2-12x-72

zepdrix (zepdrix):

\[\Large\rm f'(x)=12x^2-12x-72\]Mmm ok good. So to find increasing/decreasing we need critical points first. Those are the spots where it can change from increasing to decreasing, etc.. To find critical points we set the first derivative equal to zero and solve for x, yes?

zepdrix (zepdrix):

\[\Large\rm 0=12x^2-12x-72\]What x values do you get? :)

OpenStudy (anonymous):

so i take out 12(x^2-12x-72)

zepdrix (zepdrix):

Careful! Make sure you take a 12 out of each term, not just the first!!\[\Large\rm 0=12(x^2-x-6)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

whats next @zepdrix

zepdrix (zepdrix):

So you need to solve for x to find the critical points. Divide each side by 12,\[\Large\rm 0=x^2-x-6\]Looks like it will factor, yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x-6)(x+1)

zepdrix (zepdrix):

Hmm I don't think those work out correctly. -6 and 1 will give us -5 for our middle term. Try again silly! :U

OpenStudy (anonymous):

Oppsies (x-5)(x-1)

zepdrix (zepdrix):

?? :(

zepdrix (zepdrix):

You need a -6, So factor of -6... -6 times 1 didn't work. Other factors of -6? :o -5 and -1 are not factors.

OpenStudy (anonymous):

what?.....

zepdrix (zepdrix):

Don't remember how to factor..? :( \[\Large\rm 0=x^2-x-6\]Factor of -6 that add to -1 are going to be,\[\Large\rm -3\quad and\quad 2\]

zepdrix (zepdrix):

\[\Large\rm 0=(x-3)(x+2)\]

OpenStudy (anonymous):

it can also be -5 and -1 right?

zepdrix (zepdrix):

No. -5 times -1 gives 5. Those are NOT factors of 6.

OpenStudy (anonymous):

oh yeahhhhh....omg i cant believe i forgot

OpenStudy (anonymous):

@zepdrix the it'll be x=3 and x=2 right?

zepdrix (zepdrix):

3 and -2 yes?

OpenStudy (anonymous):

oh yes

zepdrix (zepdrix):

Ok good.. these are stationary points ( critical points ). It's locations where the function is neither increasing nor decreasing. Like the top of a hill, or the bottom of a valley. We want to know what's happening on either side of these stationary points.

OpenStudy (anonymous):

okay

zepdrix (zepdrix):

So what we'll do is, we'll plug `test points` into the first derivative, if the result is negative, the function is decreasing there. if the result is positive, the function is increasing in that location.

zepdrix (zepdrix):

Here is our derivative function with everything factored,\[\Large\rm f'(x)=(x-3)(x+2)\]

zepdrix (zepdrix):

Let's test a point on the left side of -2. How about -3?\[\Large\rm f'(-3)=12(-3-3)(-3+2)\]Does this end up being positive or negative?

zepdrix (zepdrix):

Bahhh I missed the 12 the first time I pasted the derivative function.. my bad.

OpenStudy (anonymous):

im confused

zepdrix (zepdrix):



zepdrix (zepdrix):

If the first derivative at this point is negative, then we have decreasing.

zepdrix (zepdrix):

So we plug this value (or any value smaller than -2) into our derivative function

OpenStudy (anonymous):

so we substitute x to 3 and 2 with the derivative function?

zepdrix (zepdrix):

No. x=3 and x=-2 will give us zero in the derivative function. We know that because they are critical points. We want to know what is happening `near` our critical points. That's why I chose a point near -2.

OpenStudy (anonymous):

okay so we plug in -3 with what function

OpenStudy (anonymous):

whats the derivative function again